# A Narrative Review of Identity, Data, and Location Privacy Techniques in Edge Computing and Mobile Crowdsourcing

### v2

@@ 3 Edge Computing and Privacy Concerns
Edge computing represents a paradigm shift in data processing and management, emphasizing the decentralization of computational resources to bring data storage and processing closer to the sources of data [8]. This approach is particularly beneficial for real-time and latency-sensitive applications across diverse domains such as autonomous vehicles, smart homes, emergency response systems, smart agriculture, and modernized power grids.
@@
Technological Advancements and Architectural Design: The architectural design of edge computing is inherently modular and scalable. As depicted in Figure 2, the typical architecture comprises three primary layers: the user layer, edge layer, and cloud layer. Each layer plays a distinct role in the ecosystem, with smart devices at the user layer generating data, edge servers at the edge layer processing this data in real-time, and cloud servers at the cloud layer providing long-term storage and further analytics capabilities.
@@
Integration with Smart Technologies: Edge computing is integral to the evolution of smart technologies [20]. It enables the seamless integration of IoT devices, smart sensors, and other data sources, facilitating the development of intelligent applications that can operate autonomously and interact with their environment in real-time. This integration is crucial for the advancement of smart cities, industrial automation, and personalized healthcare, among other applications.
Literature: Recent studies, such as those by [9] and [10], distinguish edge computing from cloud computing by detailing its capabilities for real-time and local data processing, highlighting emerging research, and practical applications. Research in [11] explores ’Edge AI’, examining how edge computing and AI are integrated, their applications, and the impact of AI algorithms on decision-making and data processing at the network edge. Issues of security and privacy specific to edge computing are addressed in [12], where it is noted that conventional cloud computing privacy methods are inadequate. [13] highlights concerns about location privacy in edge computing environments, while [8] reviews security and privacy preservation methods, noting a predominant focus on cloud computing over edge-specific strategies. Additionally, a survey [14] discusses edge computing’s influence on IoT development, advocating for its suitability over other computing paradigms for IoT scenarios. Furthermore, [15] presents a comprehensive review of current research on security in edge computing, pinpointing challenges in new models, application scenarios, and technological environments, with specific attention to access control, key management, privacy protection, attack mitigation, and anomaly detection.
@@ 3.1.2 Data Privacy in Edge Computing
In edge computing, ensuring data privacy is crucial due to decentralized processing and storage. Innovative approaches include partitioning deep neural networks to balance privacy with energy consumption [53], and integrating federated learning with hybrid differential privacy and adaptive compression to protect data in industrial settings [54]. Challenges in location-based services are addressed by crafting specific privacy protection strategies [25], and a novel framework for secure data sharing emphasizes privacy-preserving computations across the computing landscape [55]. Additionally, Local Differential Privacy (LDP) techniques process data while preserving personal information, though they can reduce data utility due to noise addition [56].
@@ 3.1.3 Location Privacy in Edge Computing
Ensuring location privacy within edge computing is paramount due to the proliferation of location-based services. Techniques include task offloading with reinforcement learning to protect privacy in autonomous transport systems [41], and dual K-anonymous caching strategies within edge environments to safeguard location data [57]. Dynamic strategies to protect location privacy in multi-user mobile edge systems highlight the need for agile response mechanisms in privacy protection [58]. Additional efforts focus on balancing service efficiency with privacy in location-based services [25], and task offloading frameworks that consider location privacy [58]. Vehicular crowdsourcing applications also reflect a focus on driver privacy, combining pseudonym and mix-zone techniques [59, 60].
@@
Table 3: Privacy Concerns and Solutions in Edge Computing
@@
Challenges: Despite its numerous advantages, edge computing faces several challenges, including security concerns, data privacy issues, and management complexities [18]. The distributed nature of edge computing introduces multiple points of vulnerability that must be secured against potential threats. Managing data privacy also becomes more complex as data is processed across numerous edge devices and locations. Solutions to these challenges include implementing advanced encryption techniques, establishing robust access control measures, and developing standardized protocols for data management and security.
@@ 4 Discussion
Future research should focus on enhancing the scalability and efficiency of privacy-preserving methods, addressing the challenges of de-anonymization and data utility, and exploring new cryptographic techniques to ensure robust privacy protection. The integration of artificial intelligence and machine learning into privacy protection strategies offers promising avenues for innovation. These technologies can help identify and mitigate privacy risks in real-time, providing an additional layer of security in dynamic and complex computing environments. Additionally, interdisciplinary collaboration among researchers, industry practitioners, and policymakers is essential to create standardized guidelines and frameworks for privacy protection in mobile computing.
5 Conclusion

### v3

@@ 4 Edge Computing and Privacy Concerns
Edge computing represents a paradigm shift in data processing and management. It emphasizes the decentralization of computational resources to bring data storage and processing closer to the sources of data [67]. This approach is particularly beneficial for real-time and latency-sensitive applications across diverse domains such as autonomous vehicles, smart homes, emergency response systems, smart agriculture, and modernized power grids.
@@
Technological Advancements and Architectural Design: The architectural design of edge computing is inherently modular and scalable. As depicted in Figure 6, the typical architecture comprises three primary layers: the user layer, edge layer, and cloud layer. Each layer plays a distinct role in the ecosystem, with smart devices at the user layer generating data, edge servers at the edge layer processing this data in real time, and cloud servers at the cloud layer providing long-term storage and further analytics capabilities.
@@
Integration with Smart Technologies: Edge computing is integral to the evolution of smart technologies [34]. It enables the seamless integration of IoT devices, smart sensors, and other data sources, facilitating the development of intelligent applications that can operate autonomously and interact with their environment in real time. This integration is crucial for the advancement of smart cities, industrial automation, and personalized healthcare, among other applications.
Literature: Recent studies, such as those by [6, 38], distinguish edge computing from cloud computing by detailing its capabilities for real-time and local data processing, highlighting emerging research, and practical applications. Research in [53] explores ’Edge AI’, examining how edge computing and AI are integrated, their applications, and the impact of AI algorithms on decision-making and data processing at the network edge. Issues of security and privacy specific to edge computing are addressed in [42], where it is noted that conventional cloud computing privacy methods are inadequate. Ref. [16] highlights concerns about location privacy in edge computing environments, while [67] reviews security and privacy preservation methods, noting a predominant focus on cloud computing over edge-specific strategies. Additionally, a survey [63] discusses edge computing’s influence on IoT development, advocating for its suitability over other computing paradigms for IoT scenarios. Furthermore, [66] presents a comprehensive review of current research on security in edge computing, pinpointing challenges in new models, application scenarios, and technological environments, with specific attention to access control, key management, privacy protection, attack mitigation, and anomaly detection.
@@ 4.1.2 Data Privacy in Edge Computing
In edge computing, ensuring data privacy is crucial due to decentralized processing and storage. Innovative approaches include partitioning deep neural networks to balance privacy with energy consumption [12] and integrating federated learning with hybrid differential privacy and adaptive compression to protect data in industrial settings [23]. Challenges in location-based services are addressed by crafting specific privacy protection strategies [35], and a novel framework for secure data sharing emphasizes privacy-preserving computations across the computing landscape [72]. Additionally, Local Differential Privacy (LDP) techniques process data while preserving personal information, though they can reduce data utility due to noise addition [52].
@@ 4.1.3 Location Privacy in Edge Computing
Ensuring location privacy within edge computing is paramount due to the proliferation of location-based services. Techniques include task offloading with reinforcement learning to protect privacy in autonomous transport systems [15], and dual K-anonymous caching strategies within edge environments to safeguard location data [68]. Dynamic strategies to protect location privacy in multi-user mobile edge systems highlight the need for agile response mechanisms in privacy protection [60]. Additional efforts focus on balancing service efficiency with privacy in location-based services [35] and task offloading frameworks that consider location privacy [60]. Vehicular crowdsourcing applications also reflect a focus on driver privacy, combining pseudonym and mix-zone techniques [44, 9].
@@
Challenges: Despite its numerous advantages, edge computing faces several challenges, including security concerns, data privacy issues, and management complexities [22]. The distributed nature of edge computing introduces multiple points of vulnerability that must be secured against potential threats. Managing data privacy also becomes more complex as data are processed across numerous edge devices and locations. Solutions to these challenges include implementing advanced encryption techniques, establishing robust access control measures, and developing standardized protocols for data management and security.
Mobile crowdsourcing and edge computing have revolutionized data processing and service delivery by bringing computation closer to the data source. However, these technologies pose significant privacy challenges, particularly when handling sensitive user information such as personal identity and location data. This case study examines a smart city project that integrates edge computing and mobile crowdsourcing to improve public services like traffic management, waste disposal, and public safety. Residents contribute data via mobile applications, which are processed at distributed edge devices throughout the city. This initiative faces several privacy challenges, including identity privacy, data privacy, and location privacy.
Table 3: Privacy concerns and solutions in edge computing.
@@
5 Enhancing Privacy in Mobile Crowdsourcing and Edge Computing: A Case Study
Mobile crowdsourcing and edge computing have transformed data processing and service delivery by making computation closer to the data source. However, these technologies also present privacy concerns, especially when handling sensitive user information like personal identity and location data. We consider a smart city project that integrates edge computing and mobile crowdsourcing to enhance public services such as traffic management, waste disposal, and public safety. Residents participate by sharing data via mobile applications, which are processed at edge devices distributed across the city. Traffic patterns are monitored using mobile data collected from citizens’ smartphones, while waste management facilities use data from smart bins. As residents contribute sensitive information, this project faces the following privacy challenges: identity privacy (protecting citizens’ identities), data privacy (secure processing of shared data), and location privacy (protecting individuals’ exact whereabouts).
5.1 Identity Privacy in Edge Computing
To protect identity privacy, the project implemented identity-based encryption [64] and forward security [31]. Forward security ensures that even if an encryption key is compromised, previously encrypted data remain secure. Additionally, decentralized identity management using blockchain-based [30] trust models was introduced to reduce reliance on centralized authorities. Each participant had a digital identity verified by decentralized nodes, mitigating risks of identity theft and unauthorized access.
5.2 Data Privacy with Federated Learning
For safeguarding data privacy, the project employed federated learning [70] with hybrid differential privacy. ML models are trained on data collected from citizens without centralizing their data. Instead, model updates are shared from edge devices back to a central server. This reduces the risk of data exposure while still leveraging crowdsourced data for model improvements.
5.3 Location Privacy with K-Anonymous Caching
To protect location privacy, K-anonymous caching [68]was used to obfuscate user’s exact locations by combining location data with that of other users in a given area. For instance, when a user reports parking availability, the system aggregates reports from multiple users in the vicinity to ensure individual anonymity.
5.4 Recommendations
The following recommendations are made to enhance the project’s effectiveness:
  • Establish clear guidelines for data collection, storage, and usage. Implement anonymization techniques and conduct regular audits to enhance transparency and accountability.
  • Integrate post-quantum cryptography [65] to future-proof encryption systems against quantum computing threats. Utilize privacy-preserving authentication (PPA) [33] for efficient and secure data handling.
  • Leverage blockchain for decentralized identity management to enhance trust models and improve data security with tamper-resistant record-keeping.
  • Incorporate privacy considerations from the outset of projects through thorough privacy impact assessments and embrace principles like data minimization.
  • Encourage public participation in decision-making processes related to data usage and enforce regulatory oversight to ensure compliance with evolving privacy standards.
  • Focus on scalability by integrating AI and machine learning for dynamic risk detection in real time. Implement trusted execution environments for secure processing of sensitive data.
  • Balance energy consumption with privacy needs using energy-efficient cryptographic methods that maintain security without compromising performance.
@@ 6 Discussion
Future research should focus on enhancing the scalability and efficiency of privacy-preserving methods [57, 18], addressing the challenges of de-anonymization and data utility, and exploring new cryptographic techniques to ensure robust privacy protection. The integration of artificial intelligence and machine learning into privacy protection strategies offers promising avenues for innovation. These technologies can help identify and mitigate privacy risks in real time, providing an additional layer of security in dynamic and complex computing environments. Additionally, interdisciplinary collaboration among researchers, industry practitioners, and policymakers is essential to create standardized guidelines and frameworks for privacy protection in mobile computing.
6.1 Emerging Trends in Privacy Preservation
Some key trends and emerging technologies to enhance privacy while enabling valuable insights and innovations are the following:
  1. 1.
    Artificial Intelligence (AI) and ML are increasingly being used to enhance privacy protection. These methods and technologies are already being used for anomaly detection, breach prediction, and automation of privacy-preserving processes [62].
  2. 2.
    Federated Learning is also a promising approach to privacy-preserving machine learning [24]. This technique allows models to be trained on different distributed datasets without centralizing the data and addresses privacy concerns in sectors such as healthcare and finance. The process involves the following: (1) initial model training on a central server; (2) distribution of the model to individual devices; (3) local training on each device; (4) transmission of model updates back to the central server; (5) aggregation of updates to improve the global model.
  3. 3.
    Encryption remains a main method of data protection, with many methods for enhancing its capabilities and applications. For example, homomorphic encryption [58] allows computations to be performed on encrypted data without decryption. This method allows secure data processing in untrusted environments, such as public clouds.
  4. 4.
    Post-Quantum Cryptography [65] is used for encrypting algorithms that are resistant to attacks by both classical and quantum computers. Key developments include the following: (1) NIST standardization process for post-quantum cryptographic algorithms [46]. (2) Integration of post-quantum algorithms into existing protocols [46]. (3) Hybrid approaches combining classical and post-quantum methods.
  5. 5.
    Privacy-Enhancing Computation (PEC) [47] covers a range of technologies that enable data to be processed or analyzed while keeping it encrypted or otherwise protected.
  6. 6.
    Trusted Execution Environments (TEEs) [55] provide isolated execution environments within processors and allow sensitive computations to be performed securely even on untrusted systems. This allows hardware-level isolation of sensitive code and data and protection against privileged software attacks.
  7. 7.
    Secure Multi-Party Computation (SMPC) [71] allows multiple parties to compute a function together over their inputs while keeping these inputs private. Some use cases include privacy-preserving benchmarking, secure auctions and voting systems, and machine learning in a collaborative environment across multiple data owners.
6.2 Interdisciplinary Efforts
In the domains of mobile crowdsourcing and edge computing, fostering interdisciplinary collaboration is essential to create comprehensive frameworks that balance technological advancement with user privacy. Research is also directed towards decentralized systems, which enhance user control over data and mitigate risks associated with centralized data repositories by processing data closer to its source. Blockchain technology can be integrated for traceability and immutability in data handling practices and can enhance security and transparency, allowing users to verify how their data are used and shared.
Incorporating energy efficiency into mobile crowdsourcing and edge computing is crucial for sustainable technology development [73]. One effective strategy is the use of edge computing, which enhances energy efficiency by processing data closer to where it is generated. These methods can help in reducing the need for extensive data transmission and lowering overall energy consumption. This approach can help minimize latency and improve data processing speeds and can also significantly reduce the carbon footprint associated with centralized data centers.
7 Conclusions
@@
Funding: This research received no external funding
Conflict of Interest:The authors declare no conflict of interest
Table 4: Important Definitions
Keyword Definition
Privacy The protection of sensitive information from unauthorized access, ensuring personal data such as identity, location, and activities are secure.
Security Measures taken to protect systems and data from cyber-attacks, unauthorized access, and breaches in both mobile crowdsourcing and edge computing.
Edge Computing A decentralized approach where data processing occurs closer to the data source (e.g., IoT devices) rather than relying on centralized cloud servers, enhancing performance and reducing latency.
Mobile Crowdsourcing A method where data is collected from a large number of individuals through mobile devices, often for real-time applications such as traffic monitoring or environmental sensing.
Location Privacy The protection of an individual’s geographical location information to prevent unauthorized tracking or misuse.
Data Privacy Ensuring that the collected data remains confidential and is not accessed or altered without permission, with special attention to sensitive information like health or financial data.
Identity Privacy The safeguarding of personal identifiers such as names or social security numbers to prevent identity theft or unauthorized linking of activities to an individual.
Cloud Computing The use of remote servers to store, manage, and process data, as opposed to local servers or personal devices, often integrated with edge computing for enhanced scalability.
IoT (Internet of Things) The network of connected devices, such as smartphones and smart home appliances, that collect and exchange data through the internet.
Differential Privacy A privacy technique that adds random noise to data sets, preventing the identification of individual data points while allowing analysis of aggregate trends.
Pseudonymization The process of replacing private identifiers (like names) with pseudonyms to ensure anonymity while still allowing the data to be useful.
Homomorphic Encryption A type of encryption that allows computations to be performed on encrypted data without needing to decrypt it, ensuring data privacy even during analysis.
Federated Learning A machine learning technique that allows multiple devices to collaboratively train a model without sharing raw data, preserving data privacy.
Sybil Attack A type of security attack where a single adversary creates multiple fake identities in a network to manipulate the system, commonly discussed in crowdsourcing security.
Data Poisoning A malicious attack where false or misleading data is introduced into a dataset to corrupt the outcome of a crowdsourced or machine learning model.
Spatial Cloaking A privacy technique used to blur or generalize location data to protect an individual’s exact location, commonly used in location-based services.
Blockchain A decentralized digital ledger technology that securely records transactions across multiple systems, enhancing transparency and security, especially for data privacy.
Encryption A method of converting readable data into an encoded format that can only be read by someone who has the correct decryption key, protecting data during storage and transmission.
Task Offloading The process of transferring computational tasks from resource-limited devices (e.g., mobile devices) to more powerful edge servers, improving efficiency and performance.
Trusted Execution Environment (TEE) A secure area of a device processor that ensures code and data loaded inside are protected from being tampered with the goal to provide a safe execution space for sensitive operations.